A ray of sodium light at 589nm is incident on a rectangular slab of crown glass at an angle of 45
1
Expert's answer
2021-09-21T16:03:38-0400
"n\sin{a}=\sin{45}\\a=\arcsin{\frac{\sin{45}}{1.52}}\\a=27.7\degree"